OpenStudy (anonymous):

f(x) = Square root of quantity x plus seven.; g(x) = 8x - 11 Find f(g(x)).

OpenStudy (anonymous):

\[\sqrt{x+7}\] is it tht f(x)

OpenStudy (anonymous):

yes thats f(x)

OpenStudy (anonymous):

\[\sqrt{8x-11+7} is f(g(x)) \] solve it u will get\[2\sqrt{2x+1}\]

OpenStudy (anonymous):

thank you so much!!!!